The width of a rectangle is 5cm less than half its length. Its area is 72cm^2. What is the length and width of the rectangle?
Found 2 solutions by nerdybill, sigma10:
Answer by nerdybill(7384)

The width of a rectangle is 5cm less than half its length. Its area is 72cm^2. What is the length and width of the rectangle?
.
Let x = length of rectangle
then
x/2 - 5 = width
.
x(x/2 - 5) = 72
multiplying both sides by 2:
x(x - 10) = 144
x^2 - 10x = 144
x^2 - 10x - 144 = 0
(x-18)(x+8) = 0
x = {18, -8}
Tossing out the negative solution leaves:
x = 18 cm (length)
.
width:
x/2 - 5 = 18/2 - 5 = 9 - 5 = 4 cm (width)

Answer by sigma10(1)

A = 72 cm^2 also A = w x l
w = l/2 - 5
l = l
A = (l/2 - 5)(l)= 72 solve this equation, (l-10)(l)= 2(72) and l = 18 cm
then
w = 18/2 - 5 w = 4 cm.
Answer: l=18 cm and w=4 cm